# ARCAS: An Augmented Reality Collision Avoidance System with SLAM-Based Tracking for Enhancing VRU Safety

### v1

@@ Abstract
Vulnerable road users (VRUs) face high collision risks in mixed traffic, yet most existing safety systems prioritize driver or vehicle assistance over direct VRU support. This paper presents ARCAS, a real-time augmented reality collision avoidance system that provides personalized spatial alerts to VRUs via wearable AR headsets. By fusing roadside 360-degree 3D LiDAR with SLAM-based headset tracking and an automatic 3D calibration procedure, ARCAS accurately overlays world-locked 3D bounding boxes and directional arrows onto approaching hazards in the user’s passthrough view. The system also enables multi-headset coordination through shared world anchoring. Evaluated in real-world pedestrian interactions with e-scooters and vehicles (180 trials), ARCAS nearly doubled pedestrians’ time-to-collision and increased counterparts’ reaction margins by up to 4× compared to unaided-eye conditions. Results validate the feasibility and effectiveness of LiDAR-driven AR guidance and highlight the potential of wearable AR as a promising next-generation safety tool for urban mobility.
@@ I Introduction
Traffic crashes involving vulnerable road users (VRUs) remain a major global safety concern. According to the World Health Organization, pedestrians and cyclists account for more than half of the 1.35 million annual traffic fatalities worldwide [1]. In the United States, pedestrian and cyclist deaths have increased at an alarming rate. During the first half of 2023, 3,373 pedestrians lost their lives on US streets, representing a 19% increase compared to 2019 [2]. Similarly, 1,149 cyclists were killed on US roads in 2023, marking a 4% increase from 2022. As autonomous vehicles (AVs) become more common in urban areas, the absence of a human driver necessitates alternative methods to interpret vehicle intentions to VRUs [3, 4]. Consequently, there is a growing need for systems that enhance communication and situational awareness for VRUs in mixed traffic environments.
@@
To mitigate these challenges, numerous safety technologies have been introduced, including automatic emergency braking (AEB) systems, smart crosswalks, and raised pedestrian infrastructure [5]. However, these technologies still rely on driver compliance, assuming that drivers will act logically and follow traffic laws. Prior studies have demonstrated that external human–machine interfaces (eHMIs), including LED strips, LED screens, and road projections, can compensate the lack of driver cues and support pedestrians in making more effective road-crossing decisions [6]. Despite their potential, current eHMI solutions have certain limitations, including difficulties in addressing individual pedestrians [7], challenges in visibility due to occlusions [8], and the absence of standardization [9], which can lead to inconsistent interface designs that confuse pedestrians and pose potential safety risks.
@@
A promising recent solution to these challenges involves VRU-vehicle interaction through wearable augmented reality (AR) devices. AR overlays virtual information onto the real world, allowing users to remain aware of their surroundings while receiving safety-relevant alerts [10, 9]. Unlike vehicle-mounted displays, the AR-based approach effectively enables personalized communications to individual VRUs, given that the information is no longer limited to the vehicle exterior [11, 12]. While AR has great potential for VRU safety, such applications may seem unconventional or impractical to navigate traffic safely [11]. The growing adoption of AR technologies, such as Google AR navigation on smartphones, suggests that visual assistance through AR glasses could be a feasible safety solution for VRUs in the near future.
@@
Figure 1: System overview of the proposed ARCAS. The left portion of the figure illustrates the AR headset (a Meta Quest Pro operating in passthrough mode), while the right portion shows the positioning module, consisting of a 3D 360-degree LiDAR sensor and the associated TCP server.
@@
Most existing AR studies have focused on the driver perspective, exploring head-up and windshield displays to highlight nearby pedestrians or cyclists [13, 14]. Although a growing number of studies have examined AR interfaces designed for VRUs [15], many remain conceptual or simulated. Existing solutions are designed to provide road-crossing information, such as zebra crossing and directional arrows [16, 17], and to convey collision warning signs to VRUs, including vehicle overlays [10, 12]. Only a limited number of studies have explored AR applications in real traffic environments, and these have primarily focused on evaluating interface concepts for conveying information rather than developing fully traffic collision warning system [11]. To the best of our knowledge, no prior work has demonstrated a real-time AR collision avoidance system for VRUs in real traffic or examined multi-user AR alignment for coordinated awareness.
@@
To address these research gaps, the primary objective of this paper is to develop an Augmented Reality Collision Avoidance System (ARCAS), a real-time framework designed to enhance VRU safety across dynamic traffic conditions (Figure 1). ARCAS integrates roadside 360-degree 3D LiDAR sensing with wearable AR displays to continuously track surrounding traffic, detect potential collision risks, and deliver immediate spatial alerts. A key strength of the system is its automatic 3D calibration between the LiDAR frame and the AR headset frame, enabling accurate alignment of virtual warnings with real-world objects. Additionally, ARCAS supports multi-headset operation, allowing multiple VRUs or first responders, such as firefighters, to share a unified world coordinate system for coordinated situational awareness in complex environments. More specifically, the main contributions of this novel AR-based approach are summarized as follows:
@@
    Wearable AR Visualization and Calibration: A Meta Quest Pro (MQPro) headset performs real-time functions, including the user’s 6 DoF pose using SLAM-based tracking, gaze tracking, and video streaming, while also running the calibration procedure that aligns the LiDAR coordinate frame with the headset reference frame. The headset functions as a TCP client to receive 3D target information from the server and renders spatial cues, such as bounding boxes and directional arrows, to alert users of potential collisions in real time.
@@ II Related Work
Prior research has identified human error as the leading cause of road accidents [18]. Technologies such as advanced driver assistance systems (ADAS), adaptive cruise control (ACC), and fully autonomous vehicles (FAVs) aim to improve driver and passenger safety; however, VRUs remain at high risk. For example, in interactions with FAVs, the implicit communication conveyed solely through vehicle motion may be insufficient to ensure safe VRU–vehicle interactions [17]. As a result, recent studies have focused on external communication methods and interface systems that explicitly convey a vehicle’s status and intentions to surrounding road users.
@@ II-A Vehicle-mounted Interfaces
Initial pedestrian–vehicle interface designs relied on visual signals mounted on the vehicle exterior. One early concept, “Eyes on a Car” [19], used animated eyes embedded in the headlights to mimic driver–pedestrian eye contact. The eyes tracked pedestrians to indicate yielding intent or looked forward to signal non-yielding behavior. While this design helped pedestrians make faster crossing decisions in virtual experiments, some users reported discomfort due to the Uncanny Valley effect and questioned the system’s reliability [20].
Similarly, common approach used LED strip-based interfaces installed over the windshield or along the vehicle side. These interfaces used dynamic lighting patterns to signal different vehicle states such as stopping, yielding, or resuming movement [21, 22]. Pedestrians generally found these cues intuitive, with side-mounted strips preferred for their clarity. However, some participants questioned the necessity of continuous visual feedback in all situations [20].
The “Smiling Car” interface [20] was another LED-based design that incorporated anthropomorphic features, transforming a horizontal yellow line into a smile when a pedestrian is detected to indicate yielding intent. Studies involving single pedestrian-vehicle interactions found the interface intuitive and effective in conveying clear messages [20, 23]. However, these vehicle-mounted interfaces remain limited by their reliance on visual signals, which can be difficult to perceive in poor weather, low-light conditions, or from a distance.
@@ II-B Road Projection-based Interfaces
To address the limitations of vehicle-mounted displays, researchers have investigated projection-based interfaces that project visual signals directly onto the road surface. One prototype [24] projects parallel lines in front of the vehicle, with line spacing that varies based on vehicle speed. Although this design helped pedestrians make quicker crossing decisions, their attention often shifted to the projection rather than the vehicle itself. Another approach [25] uses a color-changing projection system with three colors, red during normal driving, yellow when slowing for VRUs, and green when it is safe to cross. While these projections improve visibility and build on familiar cues such as crosswalk markings, their effectiveness can degrade under certain road conditions and may increase VRU cognitive load.
@@ II-D Wearable Augmented Reality Concepts
Wearable Augmented Reality (AR) offers many of the capabilities of smartphones while enabling more immersive interaction. Unlike traditional devices, AR overlays virtual content onto the real world, enhancing information display and situational awareness.As consumer-grade AR glasses become more accessible, VRUs may soon adopt wearable AR in everyday settings, creating new opportunities for advanced vehicle–VRU interaction paradigms [9].
Wearable AR can function as an intelligent assistant that interprets the surrounding environment, monitors user states (e.g., location, orientation, and gaze), and adapts information to situational needs. Expert interviews in recent work [9] highlight the potential of wearable AR to address scalability challenges in AV–VRU interaction. Accordingly, several studies have investigated AR-based eHMIs. Hesenius et al. [16] developed three concepts for pedestrians: one to display a visual walking path, another to highlight safe crossing zones, and a third to convey vehicle intentions and predicted stopping points. Tong and Jia [26] proposed an AR-based system that alerts pedestrians about approaching vehicles to improve situational awareness during crossings.
More recently, Tabone et al. [11] introduced a conceptual AR prototype that communicates vehicle intentions and crossing information directly to VRUs through wearable displays. In addition, Lin et al. [27] introduced an AR-based system for real-time pedestrian conflict alerts, but the system was not evaluated in real-world urban environments. While these studies offer promising concepts, most remain limited to theoretical frameworks, AR-interface designs , or single-vehicle scenarios with limited empirical validation in complex real-world environments involving multiple traffic hazards. To address these gaps, the present research aims to design an empirical, real-time AR-based collision avoidance system designed to enhance VRU safety in dynamic traffic conditions.
@@ III-A General Overview
The ARCAS system is structured around real-time communication between a roadside 3D 360-degree LiDAR sensor and a wearable MQPro headset, coordinated through a centralized TCP server running on a standard laptop. The ARCAS TCP architecture consists of a TCP client deployed on the AR headset and a corresponding TCP server interfaced with the 3D LiDAR sensor. The system components are illustrated in Figure1. Object tracking is managed by the LiDAR sensor, while all computations related to spatial mapping, calibration, and visual-cue rendering are performed locally on the AR headset. The headset software is developed in C# using the Unity engine, while the server software is implemented in C++. The software workflow operates as follows:
@@ III-B Real-Time Coordinate Transformation for AR Visualization
In this application, all users are assumed to move within a three-dimensional space defined by coordinates $(x,y,z)$. The AR headset internal tracking system, based on SLAM, provides the spatial information of the headset relative to the world frame, denoted as $(X,Y,Z)$. With the introduction of a 3D LiDAR sensor, the transformation between these reference frames is modeled as a full 3D rotation followed by a translation, as described in Equation 1.
$\left[\begin{array}[]{c}X\\ Y\\ Z\end{array}\right]=\left[\begin{array}[]{c}X_{0}\\ Y_{0}\\ Z_{0}\end{array}\right]+\mathbf{R}\left[\begin{array}[]{c}x\\ y\\ z\end{array}\right]$ (1)
After the calibration phase, the system computes the optimal values for $(X_{0},Y_{0},Z_{0})$ and the 3D rotation matrix $\mathbf{R}$. The positioning system, based on a 3D LiDAR sensor, provides real-time measurements of both the headset user position and the positions of other moving targets. The system applies the transformation defined in Equation 1 to each target to accurately render 3D bounding boxes in the headset’s world frame.
@@ III-C Kalman Filter
To reduce the measurement noise associated with the external 3D positioning system, a Kalman Filter is employed to estimate both position and velocity within the LiDAR reference frame. In the 3D configuration, the Kalman Filter is extended to estimate full three-dimensional position and velocity, resulting in a six-dimensional state vector and a 6×6 state-transition model. In contrast, the AR internal tracker introduces negligible noise and requires no additional filtering.
@@
$x_{k}=\begin{bmatrix}x(k)\\ y(k)\\ z(k)\\ v_{x}(k)\\ v_{y}(k)\\ v_{z}(k)\end{bmatrix}\quad z_{k}=\begin{bmatrix}x_{lidar}(k)\\ y_{lidar}(k)\\ z_{lidar}(k)\end{bmatrix}$
The discrete-time motion equation used for prediction is:
$x_{k+1}=Ax_{k}+w_{k}$ (2)
where $A$ is the state transition matrix and $w_{k}$ is the process noise vector.
@@
$z_{k}=Hx_{k}+v_{k}$ (3)
where $H$ is the observation matrix and $v_{k}$ is the observation noise vector:
$H=\begin{bmatrix}1&0&0&0&0&0\\ 0&1&0&0&0&0\\ 0&0&1&0&0&0\end{bmatrix}$
@@
$K_{k}$ $=P_{k|k-1}H^{\top}\left(HP_{k|k-1}H^{\top}+R\right)^{-1}$ (6)
$\hat{x}_{k|k}$ $=\hat{x}_{k|k-1}+K_{k}\left(z_{k}-H\hat{x}_{k|k-1}\right)$ (7)
$P_{k|k}$ $=\left(I-K_{k}H\right)P_{k|k-1}$ (8)
@@
$Q=\mathrm{diag}(\lambda^{2},\lambda^{2},\lambda^{2},\mu^{2},\mu^{2},\mu^{2}),\quad R=\mathrm{diag}(\sigma^{2},\sigma^{2},\sigma^{2})$

### v2

@@ Abstract
Vulnerable road users (VRUs) face high collision risks in mixed traffic, yet most existing safety systems prioritize driver or vehicle assistance over direct VRU support. This paper presents ARCAS, a real-time augmented reality (AR) collision avoidance system that provides personalized spatial alerts to VRUs via wearable AR headsets. By fusing roadside 360° 3D LiDAR with SLAM-based headset tracking and an automatic 3D calibration procedure, ARCAS accurately overlays world-locked 3D bounding boxes and directional arrows onto approaching hazards in the user’s passthrough view. The system also enables multi-headset coordination through shared world anchoring. Evaluated in real-world pedestrian interactions with e-scooters and vehicles (180 trials), ARCAS nearly doubles pedestrians’ time to collision and increases counterparts’ reaction margins by up to 4× compared to unaided eye conditions. Results validate the feasibility and effectiveness of LiDAR-driven AR guidance and highlight the potential of wearable AR as a promising next generation safety tool for urban mobility.
@@ I Introduction
Traffic crashes involving vulnerable road users (VRUs) remain a major global safety concern. According to the World Health Organization, pedestrians and cyclists account for more than half of the 1.35 million annual traffic fatalities worldwide [31]. In the United States, pedestrian and cyclist deaths have increased at an alarming rate. During the first half of 2023, 3,373 pedestrians lost their lives on US streets, representing a 19% increase compared to 2019 [27]. Similarly, 1,149 cyclists were killed on US roads in 2023, marking a 4% increase from 2022. As autonomous vehicles (AVs) become more common in urban areas, the absence of a human driver necessitates alternative methods to interpret vehicle intentions to VRUs [29, 15]. Consequently, there is a growing need for systems that enhance communication and situational awareness for VRUs in mixed traffic environments [1, 4].
@@
To mitigate these challenges, numerous safety technologies have been introduced, including automatic emergency braking systems, smart crosswalks, and raised pedestrian infrastructure [10]. However, these technologies still rely on driver compliance, assuming that drivers will act logically and follow traffic laws. Prior studies have demonstrated that external human–machine interfaces (eHMIs), including LED strips, LED screens, and road projections, can compensate the lack of driver cues and support pedestrians in making more effective road-crossing decisions [2]. Despite their potential, current eHMI solutions have certain limitations, including difficulties in addressing individual pedestrians [6], challenges in visibility due to occlusions [8], and the absence of standardization [23], which can lead to inconsistent interface designs that confuse pedestrians and pose potential safety risks.
@@
A promising recent solution to these challenges involves VRU-vehicle interaction through wearable augmented reality (AR) devices. AR overlays virtual information onto the real world, allowing users to remain aware of their surroundings while receiving safety relevant alerts [24, 23]. Unlike vehicle-mounted displays, the AR-based approach effectively enables personalized communications to individual VRUs, given that the information is no longer limited to the vehicle exterior [22, 26]. While AR has great potential for VRU safety, such applications may seem unconventional or impractical to navigate traffic safely [22]. The growing adoption of AR technologies, such as Google AR navigation on smartphones, suggests that visual assistance through AR glasses could be a feasible safety solution for VRUs in the near future.
@@
Figure 1: System overview of the proposed ARCAS. The left portion of the figure illustrates the AR headset (a Meta Quest Pro operating in passthrough mode), while the right portion shows the positioning module, consisting of a 3D 360° LiDAR sensor and the associated TCP server.
@@
Most existing AR-based studies have focused on the driver perspective, exploring head-up and windshield displays to highlight nearby pedestrians or cyclists [30]. Although a growing number of studies have examined AR interfaces designed for VRUs [18], many remain conceptual or simulated. Existing solutions are designed to provide road crossing information, such as zebra crossing and directional arrows [13, 20], and to convey collision warning signs to VRUs, including vehicle overlays [24, 26]. Only a limited number of studies have explored AR applications in real traffic environments, and these have primarily focused on evaluating interface concepts for conveying information rather than developing fully traffic collision warning system [22]. To the best of our knowledge, no prior work has demonstrated a real-time AR collision avoidance system for VRUs in real traffic or examined multi-user AR alignment for coordinated awareness.
@@
To address these research gaps, the primary objective of this paper is to develop an Augmented Reality Collision Avoidance System (ARCAS), a real-time framework designed to enhance VRU safety across dynamic traffic conditions (Figure 1). ARCAS integrates roadside 360° 3D LiDAR sensing with wearable AR displays to continuously track surrounding traffic, detect potential collision risks, and deliver immediate spatial alerts. A key strength of the system is its automatic 3D calibration between the LiDAR frame and the AR headset frame, enabling accurate alignment of virtual warnings with real-world objects. Additionally, ARCAS supports multi-headset operation, allowing multiple VRUs or first responders, such as firefighters, to share a unified world coordinate system for coordinated situational awareness in complex environments. More specifically, the main contributions of this novel AR-based approach are summarized as follows:
@@
    Wearable AR Visualization and Calibration: A Meta Quest Pro (MQPro) headset performs real-time functions, including the user’s 6 DoF pose using visual-inertial simultaneous localization and mapping (SLAM) [9], gaze tracking, and video streaming, while also running the calibration procedure that aligns the LiDAR coordinate frame with the headset reference frame. The headset functions as a TCP client to receive 3D target information from the server and renders spatial cues, such as bounding boxes and directional arrows, to alert users of potential collisions in real time.
@@ II Related Work
Prior research has identified human error as the leading cause of road accidents [12]. Technologies such as advanced driver assistance systems, adaptive cruise control, and fully AVs aim to improve driver and passenger safety [28]; however, VRUs remain at high risk. For example, in interactions with AVs, the implicit communication conveyed solely through vehicle motion may be insufficient to ensure safe VRU–vehicle interactions [20]. As a result, recent studies have focused on external communication methods and interface systems that explicitly convey a vehicle’s status and intentions to surrounding road users.
@@ II-A Vehicle-mounted Interfaces
Initial pedestrian–vehicle interface designs relied on visual signals mounted on the vehicle exterior. One early concept, “Eyes on a Car” [5], used animated eyes embedded in the headlights to mimic driver–pedestrian eye contact. The eyes tracked pedestrians to indicate yielding intent or looked forward to signal non yielding behavior. While this design helped pedestrians make faster crossing decisions in virtual experiments, some users reported discomfort due to the Uncanny Valley effect and questioned the system’s reliability [14]. Similarly, common approach used LED strip-based interfaces installed over the windshield or along the vehicle side. These interfaces used dynamic lighting patterns to signal different vehicle states such as stopping, yielding, or resuming movement [11]. Pedestrians generally found these cues intuitive, with side-mounted strips preferred for their clarity. However, some participants questioned the necessity of continuous visual feedback in all situations. The “Smiling Car” interface [17] was another LED-based design that incorporated anthropomorphic features, transforming a horizontal yellow line into a smile when a pedestrian is detected to indicate yielding intent. Studies involving single pedestrian-vehicle interactions found the interface intuitive and effective in conveying clear messages [7]. However, these vehicle-mounted interfaces remain limited by their reliance on visual signals, which can be difficult to perceive in poor weather, low light conditions, or from a distance.
@@ II-B Road Projection-based Interfaces
To address the limitations of vehicle-mounted displays, researchers have investigated projection-based interfaces that project visual signals directly onto the road surface. One prototype [3] projected parallel lines in front of the vehicle, with line spacing that varies based on vehicle speed. Although this design helped pedestrians make quicker crossing decisions, their attention often shifted to the projection rather than the vehicle itself. Another approach [19] used a color changing projection system with three colors, red during normal driving, yellow when slowing for VRUs, and green when it is safe to cross. While these projections improve visibility and build on familiar cues such as crosswalk markings, their effectiveness can degrade under certain road conditions and may increase VRU cognitive load.
@@ II-D Wearable Augmented Reality Concepts
Wearable AR offers many of the capabilities of smartphones while enabling more immersive interaction. Unlike traditional devices, AR overlays virtual content onto the real world, enhancing information display and situational awareness. As consumer grade AR glasses become more accessible, VRUs may soon adopt wearable AR in everyday settings, creating new opportunities for advanced VRU-vehicle interaction paradigms [23]. Wearable AR can function as an intelligent assistant that interprets the surrounding environment, monitors user states (e.g., location, orientation, and gaze), and adapts information to situational needs. Expert interviews in recent work [23] highlighted the potential of wearable AR to address scalability challenges in AV–VRU interaction. Accordingly, several studies have investigated AR-based eHMIs. Hesenius et al. [13] developed three concepts for pedestrians: one to display a visual walking path, another to highlight safe crossing zones, and a third to convey vehicle intentions and predicted stopping points. Tong and Jia [25] proposed an AR-based system that alerts pedestrians about approaching vehicles to improve situational awareness during crossings. More recently, Tabone et al. [22] introduced a conceptual AR prototype that communicates vehicle intentions and crossing information directly to VRUs through wearable displays. In addition, Lin et al. [16] introduced an AR-based system for real-time pedestrian conflict alerts, but the system was not evaluated in real-world urban environments. While these studies offer promising concepts, most remain limited to theoretical frameworks, AR-interface designs, or single vehicle scenarios with limited empirical validation in complex real-world environments involving multiple traffic hazards. To address these gaps, the present research aims to design an empirical, real-time AR-based collision avoidance system designed to enhance VRU safety in dynamic traffic conditions.
@@ III-A General Overview
The ARCAS system is structured around real-time communication between a roadside 3D 360° LiDAR sensor and a wearable MQPro headset, coordinated through a centralized TCP server running on a laptop (specifications detailed in Section IV-A). The ARCAS TCP architecture consists of a TCP client deployed on the AR headset and a corresponding TCP server interfaced with the 3D LiDAR sensor. The system components are illustrated in Figure 1. Object tracking is managed by the LiDAR sensor, while all computations related to spatial mapping, calibration, and visual-cue rendering are performed locally on the AR headset. The headset software is developed in C# using the Unity engine, while the server software is implemented in C++. The software workflow operates as follows:
@@ III-B Real-Time Coordinate Transformation for AR Visualization
In this application, all users are assumed to move within a 3D space defined by coordinates $(x,y,z)$. The AR headset internal tracking system, based on SLAM, provides the spatial information of the headset relative to the world frame, denoted as $(X,Y,Z)$. With the introduction of a 3D LiDAR sensor, the transformation between these reference frames is modeled as a full 3D rotation followed by a translation, as described in Equation (1).
$\left[\begin{array}[]{c}X\\ Y\\ Z\end{array}\right]=\left[\begin{array}[]{c}X_{0}\\ Y_{0}\\ Z_{0}\end{array}\right]+\mathbf{R}\left[\begin{array}[]{c}x\\ y\\ z\end{array}\right].$ (1)
After the calibration phase, the system computes the optimal values for $(X_{0},Y_{0},Z_{0})$ and the 3D rotation matrix $\mathbf{R}$. The positioning system, based on a 3D LiDAR sensor, provides real-time measurements of the positions of both headset user and other moving targets. The system applies the transformation defined in Equation (1) to each target to accurately render 3D bounding boxes in the headset’s world frame.
@@ III-C Kalman Filter
To reduce the measurement noise associated with the external 3D positioning system, a Kalman Filter is employed to estimate both position and velocity within the LiDAR reference frame. In the 3D configuration, the Kalman Filter is extended to estimate full three-dimensional position and velocity, resulting in a six-dimensional state vector and a 6×6 state transition model. In contrast, the AR internal tracker introduces negligible noise and requires no additional filtering.
@@
$x_{k}=\begin{bmatrix}x(k)\\ y(k)\\ z(k)\\ v_{x}(k)\\ v_{y}(k)\\ v_{z}(k)\end{bmatrix},\quad z_{k}=\begin{bmatrix}x_{lidar}(k)\\ y_{lidar}(k)\\ z_{lidar}(k)\end{bmatrix}$.
The discrete time motion equation used for prediction is:
$x_{k+1}=Ax_{k}+w_{k},$ (2)
where $A$ is the state transition matrix, and $w_{k}$ is the process noise vector.
@@
$z_{k}=Hx_{k}+v_{k},$ (3)
where $H$ is the observation matrix, and $v_{k}$ is the observation noise vector:
$H=\begin{bmatrix}1&0&0&0&0&0\\ 0&1&0&0&0&0\\ 0&0&1&0&0&0\end{bmatrix}$.
@@
$K_{k}$ $=P_{k|k-1}H^{\top}\left(HP_{k|k-1}H^{\top}+R\right)^{-1},$ (6)
$\hat{x}_{k|k}$ $=\hat{x}_{k|k-1}+K_{k}\left(z_{k}-H\hat{x}_{k|k-1}\right),$ (7)
$P_{k|k}$ $=\left(I-K_{k}H\right)P_{k|k-1}.$ (8)
@@
$Q=\mathrm{diag}(\lambda^{2},\lambda^{2},\lambda^{2},\mu^{2},\mu^{2},\mu^{2}),\quad R=\mathrm{diag}(\sigma^{2},\sigma^{2},\sigma^{2}).$